OpenStudy (anonymous):

Assuming that Switzerland's population is growing exponentially at a continuous rate of 0.19 percent a year and that the 1988 population was 6.9 million, write an expression for the population as a function of time in years. (Let t=0 in 1988.) P=? hint: logistic function Assuming that Switzerland's population is growing exponentially at a continuous rate of 0.19 percent a year and that the 1988 population was 6.9 million, write an expression for the population as a function of time in years. (Let t=0 in 1988.) P=? hint: logistic function @Mathematics

OpenStudy (anonymous):

dP/dt=KP P(0)=6.9x10^6 integral1/PdP=integral kdt k=.0019

OpenStudy (anonymous):

ln(P)=kt+c P=ce^kt

OpenStudy (anonymous):

you need to solve for c by setting t=o